# Hulk: A Universal Knowledge Translator for Human-Centric Tasks

### 05 Dec 2023 v2

@@ I Full Ablation results
As shown in Table I, we present the full experimental results of the ablation studies. Specifically, we explore the affects of (1) different weight sharing strategies, (2) attention mask designs in decoder, (3) task collaboration and interference, and (4) different initialization weights. One can see that the same conclusion can be observed that as more model components are shared, Hulk can achieve higher performance. Besides, omitting any task generally leads to reduced performance. Furthermore, employing different attention interaction methods for different tasks can optimize the model’s performance, compared with using pure full attention or diagnal attention. Last, using the pre-trained weights from other human-centric models (i.e., HAP [yuan2023hap] and PATH [tang2023humanbench]) can further boost our hulk’s performance. This result not only validates the robustness of Hulk but also highlight its potential for further enhancement.
@@
Moreover, in Table II, to assess the impact of the diverse and sufficient human-centric data, we train Hulk using a 60000-iteration schedule on the ablation datasets comprising 1.7M training samples, and full datasets comprising 30M training samples, respectively. Results indicate that incorporating more human-centric data and increasing the training iterations brings considerable performance improvement across all tasks.
@@
TABLE I: Complete experimental results of different ablation studies.
@@
w/o attribute 62.4 56.9 59.9 70.7 25.6 79.9 28.2 93.1 44.4 60.0 -2.1 12.2 47.9 42.6
w/o caption 62.7 57.3 59.9 70.7 25.6 80.2 79.3 77.8 92.9 46.5 61.3 -2.2 11.2 47.5 51.4
w/o skeleton 62.3 57.4 60.1 70.7 25.7 80.9 79.1 77.4 28.2 45.7 59.6 -4.8 10.0 46.7 46.6
w/o mesh&3d pose 62.6 57.4 60.3 70.8 25.5 81.0 79.8 78.0 27.9 93.1 75.8
w/o detection 63.2 57.6 60.9 71.4 26.4 79.9 77.6 27.8 93.1 46.0 59.8 -4.8 10.5 47.6 47.8
w/o 2d pose 56.6 52.3 53.0 65.0 75.3 75.1 26.6 93.5 24.5 48.4 -18.9 -3.5 37.2 39.0
w/o parsing 70.2 25.2 78.5 79.4 77.3 25.8 93.2 42.5 57.5 -2.9 10.8 46.7 40.3
@@
TABLE II: Complete experimental results of our Hulk performance on diverse human-centric tasks when scaling up data size. Experiments are both trained with a FULL 60,000-iteration schedule, showing that learn from additional data and achieve better performance on average. .
@@
II Additional Architecture Details
II-A Modality indicator
II-A1 Weight sharing.
II-A2 Shape of modality indicators.
II-B Positional Embeddings
II-B1 Positional embeddings in the encoder
Different from UniHCP [ci2023unihcp], we do not use a learnable positional embedding, instead, we use the bilinear interpolation of fixed MAE [he2022masked] positional embeddings in the encoder.
II-B2 Positional embeddings in the decoder
  • Anchor points. For the pedestrian detection task, we follow anchor DETR [wang2022anchor] to integrate the location information into modality indicators for achieving better performance. To ensure compatibility with the MAE positional embeddings, which have a shape of $14\times 14$, used in the encoded tokens part, we generate $17\times 17$ positional embeddings as fixed anchor points. This approach enables Hulk to focus on regressing the distances between these anchor points and the actual ground truth boxes. By doing so, the detection task is simplified, effectively reducing the complexity and difficulty of convergence.
  • 3D positional embeddings. For 3D pose estimation and mesh recovery tasks, Hulk needs to predict 3D digit locations using 3D positional embeddings. Following the practice in [feichtenhofer2022masked], we separate the 3D positional embeddings into a fixed $2D$ positional embedding and a fixed $1D$ positional embedding with $PE_{3D}=PE_{2D}+PE_{1D}$, introducing strong 3D location information into 3D regression tasks. The $2D$ positional embedding and $1D$ positional embedding can be generated by interpolating from fixed MAE positional embeddings.
II-C Objective function details
@@
Human Parsing. Human parsing aims to segment human parts. As human parts can be represented by their BERT features, we align the output tokens with BERT features of semantic classes using per-pixel contrastive loss. Following the practice in UniHCP [ci2023unihcp], we also supervise the global probability of a certain semantic class to enhance the performance. Given the conv-transformed output tokens $\hat{\mathbf{f}}$, we simply average pool them to represent the predicted probability of occurrence of certain semantic class $\hat{\mathbf{f}}_{avg}$. Given extracted BERT features $\mathbf{v}$ of $C$ parsing classes, , the total objective function for human parsing is as follows:
@@
3D Pose Estimation & Mesh Recovery. Following common practice [fastmetro, metro, li2022cliff], we train Hulk to learn 3D pose estimation task and mesh recovery task jointly. Similar to the pedestrian detection task, the 3D pose estimation task also has a combined objective function of semantic contrastive loss and digital regression loss while the mesh recovery task only has the digit regression loss due to undefined semantics of numerous vertices. Given predicted digital location $\hat{\mathbf{y}}^{d}$, predicted semantic features $\hat{\mathbf{f}}$ representing $C$ different joint semantics, we compute the objective function for 3D pose estimation and mesh recovery as
@@
TABLE III: Statistics of Training Datasets in the full setting.
@@
Human3.6M [h36m_pami] 62,668
LIP [lip] 30,462
Human parsing CIHP [cihp] 28,280
ModaNet [zheng2018modanet] 52,245
(7 datasets) VIP [vip] 18,469
Deep fashion [ge2019deepfashion2] 191,961
Paper Doll [paper] 1,035,825
COCO [coco] 149,813
AIC [aic] 378,352
2D pose estimation Human3.6M (pose) [h36m_pami] 312,187
Posetrack [andriluka2018posetrack] 97,174
(8 datasets) 3DPW [3dpw] 68,663
JRDB-Pose [jrdb] 139,385
MPI-INF-3DHP [mehta2017monocular] 1,031,701
AIST++ [aist] 1,015,257
RAPv2 [rapv2] 67,943
Attribute recognition PA_100k [pa] 90,000
Parse27k [PARSE27k] 27,482
(6 datasets) Market [market] 12,926
HARDHC [HARDHC] 28,336
LU-Person [fu2021unsupervised] 10,684,342
CrowdHuman [shao2018crowdhuman] 15,000
Pedestrian detection EuroCity [braun2018eurocity] 21,785
CityPersons [zhang2017citypersons] 2,778
(6 datasets) WiderPerson [zhang2019widerperson] 9,000
WiderPedestrian [loy2019wider] 58,009
COCO-person [coco] 64,115
NTU60 [shahroudy2016ntu] 40,091
Skeleton-based action NTU120 [liu2019ntu] 63,026
GYM [shao2020finegym] 2,778
(6 datasets) Diving48 [diving] 15,027
UCF101 [soomro2012ucf101] 13,320
Kinetics-400 [kay2017kinetics] 239,737
@@
CUHK-PEDS [cuhk] 68,126
(2 datasets) SYNTH-PEDES [zuo2023plip] 12,138,157
Human3.6M [h36m_pami] 312,188
3DPW [3dpw] 22,735
3D pose & Mesh COCO [coco] 40,055
MUCO [muco] 101,883
(7 datasets) UP-3D [up3d] 7,126
MPII [mpii] 14,810
GTA [gta] 1,396,913
@@
TABLE IV: Statistics of Training Datasets in the ablation setting.
@@
Human parsing Human3.6M [h36m_pami] 62,668
LIP [lip] 30,462
(3 datasets) CIHP [cihp] 28,280
@@
AIC [aic] 378,352
RAPv2 [rapv2] 67,943
Attribute recognition PA_100k [pa] 90,000
Parse27k [PARSE27k] 27,482
(5 datasets) Market [market] 12,926
HARDHC [HARDHC] 28,336
@@
CUHK-PEDS [cuhk] 68,126
(2 datasets) SYNTH-PEDES [zuo2023plip] 132,000
Human3.6M [h36m_pami] 312,188
3DPW [3dpw] 22,735
3D pose & Mesh COCO [coco] 40,055
MUCO [muco] 101,883
(6 datasets) UP-3D [up3d] 7,126
MPII [mpii] 14,810
@@
III Details of training datasets
III-A Full setting
As shown in Table III, we utilize 42 publically available datasets to form the training set for training our Hulk model, containing 30,187,836 training samples and covering eight different human-centric tasks, including human parsing, 2D pose estimation, attribute recognition, pedestrian detection, skeleton-based action recognition, image caption, 3D pose estimation, and mesh recovery.
III-B Ablation setting
In ablation stuides, to demonstrate the effectiveness of Hulk and further explore the potential of a human-centric foundation model, we conduct several ablations on a smaller training set. In Table IV, the ablation set consists of 21 datasets with about 1.7M samples, covering all eight tasks.
IV Details of dataset-wise configurations
IV-A Training
We provide detailed dataset-wise training configurations in Table V, including total batch size, batch size per GPU, the number of GPUs, and task weights. In order to better unify different tasks and different datasets, the different datasets of certain tasks (i.e., attribute recognition, pedestrian detection, skeleton-based action, image caption, 3D pose and mesh recovery) are stacked together as a integrated dataset for training, and the datasets of some other tasks (i.e., human parsinga nd 2D pose estimation) are separately trained on different GPUs with diverse task weights.
IV-B Fine-tuning
For fine-tuning porcess, we carefully tune the learning rate, total batch size, the number of iterations, and drop path rate, and report the best performances. In details, the ViT-base backbone (i.e., in Table VI) and ViT-large backbone (i.e., in Table VII) are tuned, respectively.
TABLE V: Hulk joint training setup.
@@ IV-B Fine-tuning
TABLE VI: Detailed finetuning configs with ViT-base backbone for human-centric tasks.
@@
TABLE VII: Detailed finetuning configs with ViT-large backbone for human-centric tasks.

### 27 Dec 2023 v3

@@ Appendix A Full Ablation results
As shown in Table XVI, we present the full experimental results of the ablation studies. Specifically, we explore the effects of (1) different weight-sharing strategies, (2) attention mask designs in the decoder, (3) task collaboration and interference, and (4) different initialization weights. One can see that the same conclusion can be observed that as more model components are shared, Hulk can achieve higher performance. Besides, omitting any task generally leads to reduced performance. Furthermore, employing different attention interaction methods for different tasks can optimize the model’s performance, compared with using pure full attention or diagonal attention. Last, using the pre-trained weights from other human-centric models (i.e., HAP [31] and PATH [30]) can further boost our hulk’s performance. This result not only validates the robustness of Hulk but also highlights its potential for further enhancement.
@@
Moreover, in Table XVII, to assess the impact of the diverse and sufficient human-centric data, we train Hulk using a 60000-iteration schedule on the ablation datasets comprising 1.7M training samples, and full datasets comprising 30M training samples, respectively. Results indicate that incorporating more human-centric data and increasing the training iterations brings considerable performance improvement across all tasks.
Appendix B Additional Architecture Details
B-A Modality indicator
B-A1 Weight sharing.
B-A2 Shape of modality indicators.
B-B Positional Embeddings
B-B1 Positional embeddings in the encoder
Different from UniHCP [2], we do not use a learnable positional embedding, instead, we use the bilinear interpolation of fixed MAE [99] positional embeddings in the encoder.
B-B2 Positional embeddings in the decoder
  • Anchor points. For the pedestrian detection task, we follow anchor DETR [146] to integrate the location information into modality indicators for achieving better performance. To ensure compatibility with the MAE positional embeddings, which have a shape of $14\times 14$, used in the encoded tokens part, we generate $17\times 17$ positional embeddings as fixed anchor points. This approach enables Hulk to focus on regressing the distances between these anchor points and the actual ground truth boxes. By doing so, the detection task is simplified, effectively reducing the complexity and difficulty of convergence.
  • 3D positional embeddings. For 3D pose estimation and mesh recovery tasks, Hulk needs to predict 3D digit locations using 3D positional embeddings. Following the practice in [147], we separate the 3D positional embeddings into a fixed $2D$ positional embedding and a fixed $1D$ positional embedding with $PE_{3D}=PE_{2D}+PE_{1D}$, introducing strong 3D location information into 3D regression tasks. The $2D$ positional embedding and $1D$ positional embedding can be generated by interpolating from fixed MAE positional embeddings.
@@ B-B2 Positional embeddings in the decoder
TABLE XVI: Complete experimental results of different ablation studies. The average performances in the ablation of task collaboration and interference are not computed and replaced by N/A.
@@
w/o attribute 62.4 56.9 59.9 70.7 25.6 79.9 - - 28.2 93.1 44.4 60.0 -2.1 12.2 47.9 N/A
w/o caption 62.7 57.3 59.9 70.7 25.6 80.2 79.3 77.8 - 92.9 46.5 61.3 -2.2 11.2 47.5 N/A
w/o skeleton 62.3 57.4 60.1 70.7 25.7 80.9 79.1 77.4 28.2 - 45.7 59.6 -4.8 10.0 46.7 N/A
w/o mesh&3d pose 62.6 57.4 60.3 70.8 25.5 81.0 79.8 78.0 27.9 93.1 - - - - - N/A
w/o detection 63.2 57.6 60.9 71.4 26.4 - 79.9 77.6 27.8 93.1 46.0 59.8 -4.8 10.5 47.6 N/A
w/o 2d pose 56.6 52.3 53.0 - - 65.0 75.3 75.1 26.6 93.5 24.5 48.4 -18.9 -3.5 37.2 N/A
w/o parsing - - - 70.2 25.2 78.5 79.4 77.3 25.8 93.2 42.5 57.5 -2.9 10.8 46.7 N/A
@@
TABLE XVII: Complete experimental results of our Hulk performance on diverse human-centric tasks when scaling up data size. Experiments are both trained with a FULL 60,000-iteration schedule, showing that learn from additional data and achieve better performance on average.
@@
B-C Objective function details
@@
Human Parsing. Human parsing aims to segment human parts. As human parts can be represented by their BERT features, we align the output tokens with BERT features of semantic classes using per-pixel contrastive loss. Following the practice in UniHCP [2], we also supervise the global probability of a certain semantic class to enhance the performance. Given the conv-transformed output tokens $\hat{\mathbf{f}}$, we simply average pool them to represent the predicted probability of occurrence of certain semantic class $\hat{\mathbf{f}}_{avg}$. Given extracted BERT features $\mathbf{v}$ of $C$ parsing classes, , the total objective function for human parsing is as follows:
@@
3D Pose Estimation & Mesh Recovery. Following common practice [40, 39, 133], we train Hulk to learn 3D pose estimation task and mesh recovery task jointly. Similar to the pedestrian detection task, the 3D pose estimation task also has a combined objective function of semantic contrastive loss and digital regression loss while the mesh recovery task only has the digit regression loss due to undefined semantics of numerous vertices. Given predicted digital location $\hat{\mathbf{y}}^{d}$, predicted semantic features $\hat{\mathbf{f}}$ representing $C$ different joint semantics, we compute the objective function for 3D pose estimation and mesh recovery as
@@
TABLE XVIII: Statistics of Training Datasets in the full setting.
@@
Human3.6M [71] 62,668
LIP [75] 30,462
Human parsing CIHP [74] 28,280
ModaNet [148] 52,245
(7 datasets) VIP [149] 18,469
Deep fashion [150] 191,961
Paper Doll [151] 1,035,825
COCO [72] 149,813
AIC [73] 378,352
2D pose estimation Human3.6M (pose) [71] 312,187
Posetrack [83] 97,174
(8 datasets) 3DPW [70] 68,663
JRDB-Pose [152] 139,385
MPI-INF-3DHP [153] 1,031,701
AIST++ [154] 1,015,257
RAPv2 [76] 67,943
Attribute recognition PA_100k [77] 90,000
Parse27k [84] 27,482
(6 datasets) Market [155] 12,926
HARDHC [156] 28,336
LU-Person [6] 10,684,342
CrowdHuman [78] 15,000
Pedestrian detection EuroCity [85] 21,785
CityPersons [157] 2,778
(6 datasets) WiderPerson [86] 9,000
WiderPedestrian [158] 58,009
COCO-person [72] 64,115
NTU60 [87] 40,091
Skeleton-based action NTU120 [114] 63,026
GYM [88] 2,778
(6 datasets) Diving48 [89] 15,027
UCF101 [159] 13,320
Kinetics-400 [160] 239,737
@@
CUHK-PEDES [90] 68,126
(2 datasets) SYNTH-PEDES [79] 12,138,157
Human3.6M [71] 312,188
3DPW [70] 22,735
3D pose & Mesh COCO [72] 40,055
MUCO [91] 101,883
(7 datasets) UP-3D [161] 7,126
MPII [162] 14,810
GTA [92] 1,396,913
@@
TABLE XIX: Statistics of Training Datasets in the ablation setting.
@@
Human parsing Human3.6M [71] 62,668
LIP [75] 30,462
(3 datasets) CIHP [74] 28,280
@@
AIC [73] 378,352
RAPv2 [76] 67,943
Attribute recognition PA-100k [77] 90,000
Parse27k [84] 27,482
(5 datasets) Market [155] 12,926
HARDHC [156] 28,336
@@
CUHK-PEDES [90] 68,126
(2 datasets) SYNTH-PEDES [79] 132,000
Human3.6M [71] 312,188
3DPW [70] 22,735
3D pose & Mesh COCO [72] 40,055
MUCO [91] 101,883
(6 datasets) UP-3D [161] 7,126
MPII [162] 14,810
@@
Appendix C Potential Negative Impacts
As a foundation model for human-centric tasks, Hulk’s extensive data requirements and prolonged training time raise concerns about environmental impact due to high energy consumption. To mitigate this, future research should focus on improving computational efficiency. Additionally, Hulk is trained and evaluated on several human-centric datasets, thus may contain biases in these datasets, which need continuous effort to address to ensure ethical application.
Appendix D Details of training datasets
D-A Full setting
As shown in Table XVIII, we utilize 42 publically available datasets to form the training set for training our Hulk model, containing 30,187,836 training samples and covering eight different human-centric tasks, including human parsing, 2D pose estimation, attribute recognition, pedestrian detection, skeleton-based action recognition, image caption, 3D pose estimation, and mesh recovery.
D-B Ablation setting
In ablation studies, to demonstrate the effectiveness of Hulk and further explore the potential of a human-centric foundation model, we conduct several ablations on a smaller training set. In Table XIX, the ablation set consists of 21 datasets with about 1.7M samples, covering all eight tasks.
Appendix E Details of dataset-wise configurations
E-A Training
We provide detailed dataset-wise training configurations in Table XX, including total batch size, batch size per GPU, the number of GPUs, and task weights. In order to better unify different tasks and different datasets, the different datasets of certain tasks (i.e., attribute recognition, pedestrian detection, skeleton-based action, image caption, 3D pose, and mesh recovery) are stacked together as an integrated dataset for training, and the datasets of some other tasks (i.e., human parsinga nd 2D pose estimation) are separately trained on different GPUs with diverse task weights.
E-B Fine-tuning
For the fine-tuning process, we carefully tune the learning rate, total batch size, the number of iterations, and drop path rate, and report the best performances. In detail, the ViT-base backbone (i.e., in Table XXI) and ViT-large backbone (i.e., in Table XXII) are tuned, respectively.
TABLE XX: Hulk joint training setup.
@@ E-B Fine-tuning
TABLE XXI: Detailed finetuning configs with ViT-base backbone for human-centric tasks.
@@
TABLE XXII: Detailed finetuning configs with ViT-large backbone for human-centric tasks.